# The Beilinson–Bloch conjecture for some non-isotrivial varieties over global function fields

### v2

@@ 3. Motives
In this section, we recall some needed background on motives, and prove a motivic relation between a smooth cubic threefold and its intermediate Jacobian (proposition 3.4). Some of our arguments will rest on the theory of finite-dimensional motives [Kimura2005, Definition 3.7], in which the nilpotence theorem plays a fundamental role.
Proposition 3.1 ([Kimura2005, Proposition 7.5]).
@@
If $M$ is a motive over $k$ and $L/k$ a field extension, then $M_{L}$ is finite-dimensional iff $M$ is [Vial2017, Theorem 2].
A motive of abelian type is an object of the thick and rigid subcategory of $\mathcal{M}(k)$ generated by the motives of abelian varieties. For instance, the motive of a smooth, projective, geometrically connected curve is of abelian type [Scholl1994, Proposition 4.5]. When $m\geq 1$ and $\operatorname{char}(k)\nmid m$, the Fermat hypersurface $X\subset\mathbb{P}^{n+1}_{k}$ defined by the equation
@@
has motive of abelian type [KS79, Theorem 1]. Motives of abelian type are finite-dimensional [Kimura2005, Example 9.1].
Let $X$ be a smooth projective scheme over $k$ of pure dimension $d$, and let $H$ be a classical Weil cohomology theory on $\mathcal{M}(k)$ [Andre2004, 3.4].
@@
A Chow–Künneth decomposition of the motive of $X$ (with respect to $H$) consists of an isomorphism of motives
$\mathfrak{h}(X)\cong\bigoplus\limits_{0\leq i\leq 2d}\mathfrak{h}^{i}(X),$
such that for all $0\leq i\leq 2d$ and $j\in{\mathbb{Z}}$, $H^{j}(\mathfrak{h}^{i}(X))=0$ whenever $i\neq j$. A Chow–Künneth decomposition is said to be self-dual if for all $i$, the canonical isomorphism $\mathfrak{h}(X)^{\vee}\cong\mathfrak{h}(X)(d)$ induces an isomorphism $\mathfrak{h}^{i}(X)^{\vee}\cong\mathfrak{h}^{2d-i}(X)(d)$. Equivalently, if $\pi_{i}:\mathfrak{h}(X)\to\mathfrak{h}^{i}(X)$ denotes the projector onto the $i$-th Chow–Künneth component, the decomposition is self-dual iff for all $i$, $\pi_{i}$ is the transpose of $\pi_{2d-i}$.
By [Andre2004, Théorème 4.2.5.2 and Remarque 5.1.1.2], the definition does not depend on the choice of $H$, so we will simply speak of Chow–Künneth decompositions, with no reference to $H$. Self-dual Chow–Künneth decompositions are known to exist for motives of abelian varieties [DM1991], curves, surfaces [Murre1990], complete intersections in $\mathbb{P}^{n}$ [KS2016], and several other classes of varieties. All the concrete examples of Chow–Künneth decompositions considered in this paper will have the property that $\operatorname{CH}^{i}(\mathfrak{h}^{1}(X))=0$ when $i\neq 1$, and $\operatorname{CH}^{1}(\mathfrak{h}^{1}(X))=\operatorname{CH}^{1}(X)_{0}$ (see [Scholl1994, Theorem 4.4]).
@@
A smooth complete intersection $X\subset\mathbb{P}^{r}_{k}$ of dimension $d$ has a Chow–Künneth decomposition of the form
@@
The existence and form of the Chow–Künneth decomposition is proved in [KS2016, Section 3.5]. Suppose $X$ is a smooth cubic hypersurface of dimension $\leq 3$. Curves have finite-dimensional motive, so we may assume $\dim X\geq 2$. When $X$ has a rational point, $X$ is unirational by [Kollar2002, Theorem 1], and [Vial2017, Theorem 3.12 and Proposition 3.13] then imply that $\mathfrak{h}(X)$ is finite-dimensional. Theorem 2 of loc. cit. yields the same conclusion when $X$ does not have a rational point. ∎
The following proposition is likely known to experts, but we did not find a proof over an arbitrary field in the literature. It was proved by Shermenev over an algebraically closed field [Shermenev1970], via a different method. See [Huybrechts2023, Remark 5.3.17] for some relevant background and references.
@@
Let $X$ be a smooth cubic threefold over $k$, equipped with a self-dual Chow–Künneth decomposition. There is a canonical isomorphism
@@
as in [Scholl1994, Theorem 4.4], so that $\mathfrak{h}^{1}(F)\cong\mathfrak{h}^{1}(A)$. Let $I$ be the incidence correspondence

### v3

@@ 3. Motives
In this section, we recall some needed background on motives, and prove a motivic relation between a smooth cubic threefold and its intermediate Jacobian (proposition 3.12). An analogous relation holds for a smooth cubic fivefold, and for certain other complete intersections (corollary 3.21). The proofs of the latter results rely on standard specialization techniques for Chow motives, which enable the generalization of certain statements about motives over $\mathbb{C}$ to more-or-less arbitrary base fields.
The section is organized as follows. In subsections 3.1-3.4, we review some known facts about motives, including aspects of finite-dimensional motives and Chow–Künneth decompositions. In subsection 3.5 we give an explicit description of the motive of a smooth cubic threefold, in terms of that of its intermediate Jacobian. In subsection 3.6, we note that Fulton’s specialization map on Chow groups [38, Chapter 20.3] can be extended to a specialization functor on motives, and record some properties of this functor. In subsection 3.7, we use specialization to generalize results of Vial [99] and Achter–Casalaina-Martin–Vial [2], on representability of Chow groups and motives over characteristic zero fields, to arbitrary base fields. In subsection 3.8, we apply this to explicit examples, constructing “intermediate Jacobians” (up to isogeny) of smooth cubic fivefolds, and of some other smooth complete intersections. Though it is not needed for our immediate applications in this paper, we also generalize to essentially arbitrary base fields other examples of varieties with finite-dimensional motive and motive of abelian type, originally due to Diaz [31], Laterveer ([63], [64], [65], [66]), Bolognesi–Laterveer [11], and Fu–Moonen [35]. These examples include the Fano varieties of lines on smooth cubic threefolds and fivefolds, Gushel–Mukai threefolds and fivefolds, and a few explicit families of K3 surfaces and cubic fourfolds.
3.1. Finite-dimensional motives
Our later arguments will make ample use of the theory of finite-dimensional motives [60, Definition 3.7], in which the below result, known as the nilpotence theorem, plays a fundamental role.
Proposition 3.1 ([60, Proposition 7.5]).
@@ 3.1. Finite-dimensional motives
We will repeatedly refer to the fact, due to Vial, that if $M$ is a motive over $k$, and $L/k$ is a field extension, then $M_{L}$ is finite-dimensional iff $M$ is finite-dimensional [100, Theorem 2].
The following standard lemma is quoted from [100, Section 3.3]. It implies that two finite-dimensional motives which are isomorphic modulo numerical equivalence are in fact isomorphic.
Lemma 3.2.
Let $f:M\to N$ be a map of motives over $k$, and $f_{\operatorname{num}}:M_{\operatorname{num}}\to N_{\operatorname{num}}$ the induced map modulo numerical equivalence. If $N$ is finite-dimensional and $f_{\operatorname{num}}$ has a right inverse, then $f$ has a right inverse. If $M$ is finite-dimensional and $f_{\operatorname{num}}$ has a left inverse, then $f$ has a left inverse.
Proof.
Suppose that $N$ is finite-dimensional, and let $g_{\operatorname{num}}$ be a right inverse of $f_{\operatorname{num}}$. Let $g:N\to M$ be a lift of $g_{\operatorname{num}}$ modulo rational equivalence. Then by proposition 3.1, there exists $n\geq 0$ such that $(\operatorname{id}_{N}-f\circ g)^{n}=0$. By algebra we obtain an explicit $g^{\prime}:N\to M$ with $\operatorname{id}_{N}=f\circ g^{\prime}$. The proof of the opposite assertion is identical. ∎
In view of the previous lemma, it will be useful to have a cohomological criterion for isomorphism of numerical motives.
Lemma 3.3.
Let $M$, NN be motives over $k$, and let $f:M\to N$ and $g:N\to M$ be two maps which induce isomorphisms on $\ell$-adic cohomology. Then $f$ and $g$ induce isomorphisms modulo numerical equivalence.
Proof.
[…]
Combining lemmas 3.2 and 3.3 yields the following corollary.
Corollary 3.4.
With hypotheses and notation as in lemma 3.3, if $M$ and $N$ are further assumed to be finite-dimensional, then $M\cong N$.
3.2. Artin motives and motives of abelian type
An Artin motive is a motive which is isomorphic to a direct summand of $\mathfrak{h}(\operatorname{Spec}A)$, where $A$ is a finite étale $k$-algebra. The functor $M\mapsto\operatorname{CH}^{0}(M_{\overline{k}})$ is an equivalence from the category of Artin motives over $k$ to the category of continuous $G_{k}$-representations on finite-dimensional $\mathbb{Q}$}-vector spaces [5, Exemples 4.1.6.1]. Here the vector spaces are equipped with the discrete topology.
A motive of abelian type is an object of the thick and rigid tensor subcategory of $\mathcal{M}(k)$ generated by the motives of abelian varieties. For instance, the motive of a smooth, projective, geometrically connected curve is of abelian type [87, Proposition 4.5]. When $m\geq 1$ and $\operatorname{char}(k)\nmid m$, the Fermat hypersurface $X\subset\mathbb{P}^{n+1}_{k}$ defined by the equation
@@ 3.2. Artin motives and motives of abelian type
has motive of abelian type [58, Theorem 1]. Artin motives, and motives of abelian type, are finite-dimensional [60, Example 9.1].
The property of being an Artin motive descends along field extensions.
Lemma 3.5.
If $M\in\mathcal{M}(k)$, and there exists a field extension $L/k$ such that $M_{L}$ is an Artin motive, then $M$ is an Artin motive.
Proof.
We may assume that $L$ is algebraically closed, so that $M_{L}\cong\mathbf{1}_{L}^{\oplus r}$ for some $r\geq 0$. Then $M$ is finite-dimensional by [100, Theorem 2]. Combining lemma 3.2 with [55, Proposition 5.5] shows that $M_{\overline{k}}\cong\mathbf{1}_{\overline{k}}^{\oplus r}$. The finite-dimensional $G_{k}$-representation $\operatorname{CH}^{0}(M_{\overline{k}})$ then has a corresponding Artin motive $N\in\mathcal{M}(k)$. Since $\operatorname{CH}^{0}(M_{\overline{k}})=\operatorname{Hom}_{\mathcal{M}(\overline{k})}(\mathbf{1}_{\overline{k}},M_{\overline{k}})$, there is a $G_{k}$-equivariant isomorphism $f_{\overline{k}}:N_{\overline{k}}\xrightarrow{\sim}M_{\overline{k}}$. By [14, Lemma 2.2], this map descends to an isomorphism $f:N\xrightarrow{\sim}M$ defined over $k$. ∎
3.3. Chow–Künneth decompositions
Let $H$ be a classical Weil cohomology theory on $\mathcal{M}(k)$ [5, 3.4]; for instance, $H$ can be étale cohomology with coefficients in $\mathbb{Q}$$_{\ell}$.
@@ 3.3. Chow–Künneth decompositions
A Chow–Künneth decomposition of a motive $M$ (with respect to $H$) consists of an isomorphism of motives
$M\cong\bigoplus\limits_{i\in{\mathbb{Z}}}\mathfrak{h}^{i}(M),$
such that for all $i,j\in\mathbb{Z}$, we have $H^{j}(\mathfrak{h}^{i}(M))=0$ whenever $i\neq j$.
If $X$ is a smooth projective scheme over $k$, then we denote the components of a Chow–Künneth decomposition of $\mathfrak{h}(X)$ by $\mathfrak{h}^{i}(X)$. Suppose that $X$ has pure dimension $d$. A Chow–Künneth decomposition of $\mathfrak{h}(X)$ is then said to be self-dual if for all $i$, the canonical isomorphism $\mathfrak{h}(X)^{\vee}\cong\mathfrak{h}(X)(d)$ induces an isomorphism $\mathfrak{h}^{i}(X)^{\vee}\cong\mathfrak{h}^{2d-i}(X)(d)$. Equivalently, if $\pi_{i}:\mathfrak{h}(X)\to\mathfrak{h}^{i}(X)$ denotes the projector onto the $i$-th Chow–Künneth component, the decomposition is self-dual iff for all $i$, $\pi_{i}$ is the transpose of $\pi_{2d-i}$.
By [5, Théorème 4.2.5.2 and Remarque 5.1.1.2], the above definition does not depend on the choice of $H$, so we will simply speak of Chow–Künneth decompositions, with no reference to $H$. Self-dual Chow–Künneth decompositions are known to exist for motives of abelian varieties [30], curves, surfaces [82], complete intersections in $\mathbb{P}^{n}$ [53], and several other classes of varieties. All the concrete examples of Chow–Künneth decompositions considered in this paper will have the property that $\operatorname{CH}^{i}(\mathfrak{h}^{1}(X))=0$ when $i\neq 1$, and $\operatorname{CH}^{1}(\mathfrak{h}^{1}(X))=\operatorname{CH}^{1}(X)_{0}$ (see [87, Theorem 4.4]).
For later use, we recall the explicit Chow–Künneth decompositions of smooth complete intersections.
@@
A smooth complete intersection $X\subset\mathbb{P}^{r}_{k}$ of dimension $d$ has a self-dual Chow–Künneth decomposition of the form
@@
Proof.
The existence and form of the Chow–Künneth decomposition is proved in [53, Section 3.5]. ∎
Up to isomorphism, Chow–Künneth components of finite-dimensional motives are independent of the choice of decomposition.
Lemma 3.8.
If $M$ is a finite-dimensional motive, then for any two Chow–Künneth decompositions $\mathfrak{h}^{\bullet}(M),\mathfrak{h}^{\bullet}(M)^{\prime}$ of $M$, we have $\mathfrak{h}^{i}(M)\cong\mathfrak{h}^{i}(M)^{\prime}$ for all $i$.
Proof.
Using the Chow–Künneth projectors to construct a map $\mathfrak{h}^{i}(M)\to\mathfrak{h}^{i}(M)^{\prime}$, this follows immediately from corollary 3.4. ∎
We note that all maps between $\mathfrak{h}^{1}$s of abelian varieties arise from maps in the isogeny category of abelian varieties.
Lemma 3.9.
For $A,B$ abelian varieties over $k$, there is a canonical isomorphism
$\operatorname{Hom}_{\mathcal{M}(k)}(\mathfrak{h}^{1}(A),\mathfrak{h}^{1}(B))\cong\operatorname{Hom}_{k}(A^{\vee},B^{\vee})\otimes_{\mathbb{Z}}{\mathbb{Q}},$
where $A^{\vee}$ and $B^{\vee}$ denote the dual abelian varieties of $A$ and $B$, and $\operatorname{Hom}_{k}(A^{\vee},B^{\vee})$ denotes homomorphisms of abelian varieties. In particular, if a map $f:\mathfrak{h}^{1}(A)\to\mathfrak{h}^{1}(B)$ induces an isomorphism on $\ell$-adic cohomology, then some integer multiple of $f$ is induced by an isogeny from $B$ to $A$, so that $f$ is an isomorphism of motives.
Proof.
This follows from [87, Proposition 4.5]. ∎
3.4. The blowup formula
We will later need the below blowup formula for motives.
Proposition 3.10 ([43, Corollary 3.5]).
If $X$ is a smooth projective variety over $k$, $Z\subseteq X$ is a smooth closed subvariety of codimension $c$, and $\widetilde{X}$ is the blowup of $X$ along $Z$, then there is an isomorphism
for all $i$.
3.5. Cubic threefolds
In proposition 3.12 below, we explicitly describe the motives of smooth cubic threefolds.
Lemma 3.11.
@@ 3.5. Cubic threefolds
Curves have finite-dimensional motive, so we may assume $\dim X\geq 2$. When $X$ has a rational point, $X$ is unirational by [61, Theorem 1], and [100, Theorem 3.12 and Proposition 3.13] then imply that $\mathfrak{h}(X)$ is finite-dimensional. Theorem 2 of loc. cit. yields the same conclusion when $X$ does not have a rational point. ∎
The following proposition is likely known to experts, but we did not find a proof over an arbitrary field in the literature. It was proved by Shermenev over an algebraically closed field [88], via a different method. See [46, Remark 5.3.17] for some relevant background and references.
@@
Let $X$ be a smooth cubic threefold over $k$, equipped with a Chow–Künneth decomposition. There is then an isomorphism
@@
as in [87, Theorem 4.4], so that $\mathfrak{h}^{1}(F)\cong\mathfrak{h}^{1}(A)$. Let $I$ be the incidence correspondence